### v1

# FedADMM-InSa: An Inexact and Self-Adaptive ADMM for Federated Learning

Abstract. Federated learning (FL) is a promising framework for learning from distributed data while maintaining privacy. The development of efficient FL algorithms encounters various challenges, including heterogeneous data and systems, limited communication capacities, and constrained local computational resources. Recently developed FedADMM methods show great resilience to both data and system heterogeneity. However, they still suffer from performance deterioration if the hyperparameters are not carefully tuned. To address this issue, we propose an inexact and self-adaptive FedADMM algorithm, termed FedADMM-InSa. First, we design an inexactness criterion for the clients’ local updates to eliminate the need for empirically setting the local training accuracy. This inexactness criterion can be assessed by each client independently based on its unique condition, thereby reducing the local computational cost and mitigating the undesirable straggle effect. The convergence of the resulting inexact ADMM is proved under the assumption of strongly convex loss functions. Additionally, we present a self-adaptive scheme that dynamically adjusts each client’s penalty parameter, enhancing algorithm robustness by mitigating the need for empirical penalty parameter choices for each client. Extensive numerical experiments on both synthetic and real-world datasets are conducted. As validated by some numerical tests, our proposed algorithm can reduce the clients’ local computational load significantly and also accelerate the learning process compared to the vanilla FedADMM.
1 Introduction
With the increasing volume of data generated by massively distributed devices and organizations, traditional centralized deep learning paradigms encounter challenges in terms of data collection, privacy concerns, and scalability. To address the above challenges, federated learning (FL) [21, 22, 26] has emerged as a promising paradigm and has gained significant attention in recent years. In FL, multiple clients (devices) in a distributed environment collaboratively train a neural network model under the coordination of a central server, without centralizing their data. The training process of FL consists of several communication rounds. For example, in the most commonly used FedAvg [26] algorithm, at every communication round, each client computes a model update on its local data using the latest copy of the global model parameters and then sends the model update to the central server. The server aggregates these updates by averaging to update the global model parameters, which are then sent to all clients. Despite the fact that no raw data is transmitted, there is still a risk of privacy breaches, such as data reconstruction attacks [7, 34, 35, 37]. One of the key properties of FL is that the clients are massively distributed and have limited communication capacity [18, 26, 40]. To decrease the number of communication rounds needed to train a model, FedAvg suggests increasing local computation on each client. Unlike traditional distributed optimization [43] where consensus is performed after every local gradient computation, in FedAvg, the clients perform multiple full-batch or stochastic gradient descent steps before these local model parameters are aggregated in order to update the global model parameters. However, simply increasing local gradient descent steps as in FedAvg would cause the client drift [19] issue due to the inherent data and system heterogeneity in FL [18, 26]. The client drift issue comes from the fact that clients in FL normally possess unbalanced and non-independent and identically distributed (non-IID) local datasets, which leads to inconsistent minima of the local and global objective functions [19, 32]. As a result, the aggregated global model suffers from deviation and does not converge to its true optimum [24]. To address this, FedProx [23] was proposed by adding an extra proximal term to the client’s loss function to constrain the client’s local model to be close to the server’s global model. However, the extra proximal term might degrade the training performance unless carefully tuned [31]. In [19], a method called Scaffold is proposed to reduce client drift by using variance reduction to correct each client’s local updates. However, this method doubles the size of variables that need to be communicated between the server and clients, which is not ideal given the limited communication resources in FL. Recent findings show that FedADMM methods [1, 13, 33, 41, 42] based on the alternating direction method of multipliers (ADMM) [4, 9] are inherently resilient to heterogeneous clients in FL. FedADMM leverages dual variables to tackle statistical heterogeneity and accommodates system heterogeneity by tolerating variable amounts of work performed by clients. FedADMM maintains identical communication costs per round as FedAvg/Prox and generalizes them using the augmented Lagrangian with dual variables and an extra quadratic penalty term as the client’s local loss function. This modification strikes a balance between updating the client’s local model and staying consistent with the global model [13]. Despite the above-mentioned nice features, current FedADMM methods still suffer from performance deterioration and the straggler effect [18, 30] if the hyperparameters are not carefully tuned, especially the amount of local training workload and the choice of the penalty parameter. In particular, FedADMM methods require clients to perform a certain amount of local training workload, which corresponds to solving the ADMM subproblem inexactly. This is implemented empirically by either commanding clients to solve the subproblem to a constant given accuracy [13, 41, 42] or performing a fixed number of (stochastic) gradient descent steps [1, 33]. However, the accuracy of this solution significantly impacts the effectiveness of the algorithm [11]. Meanwhile, empirically assigning the same amount of local training workload overlooks the heterogeneity in clients’ data and systems (e.g., non-IID datasets and varying computational resources). This oversight may also cause a severe straggler effect, as waiting for the resource-constrained clients to finish the overload work will slow down the training process, and simply dropping them will lead to a biased global model [3, 42]. Another critical issue involves the selection of the penalty parameter of the quadratic term in the augmented Lagrangian functions. It has been shown in ADMM applications [14, 28, 36] that the efficiency of ADMM heavily depends on the penalty parameter. If the penalty parameter is chosen too small or too large, the solution time can increase significantly. The problem is more complicated in FedADMM as each client can use a different penalty parameter, and an inappropriate choice of the penalty parameter can potentially deteriorate the performance of FedADMM methods. To address the aforementioned issues, we propose an inexact and self-adaptive FedADMM algorithm, referred to as FedADMM-InSa. Firstly, we introduce an easy-to-implement and adaptive inexactness criterion to guide the client’s local training. Our approach eliminates the need to manually set gradient descent steps or predefine a constant accuracy. This provides each client with the flexibility to solve its subproblem inexactly based on its unique situation, thereby eliminating the potential straggler effect. Furthermore, we design a scheme to dynamically adjust each client’s penalty parameter based on the discrepancy between its local model parameters and the global model parameters, avoiding unexpected performance deterioration due to improperly chosen fixed penalty parameters. Overall, our main contributions are as follows:
  1. 1.
    We propose an inexactness criterion that enables each client to dynamically adjust the precision of local training in each communication round. There is no need to empirically set the number of gradient descent steps or perceived constant accuracy a priori. This flexibility allows our algorithm to better adapt to the heterogeneous clients and datasets, save local computational resources, and mitigate the straggler effect.
  2. 2.
    We develop a self-adaptive scheme for adjusting each client’s penalty parameter. The scheme dynamically balances the primal and dual residuals defined by the dissimilarity between the client’s local parameters and the server’s global parameters between two communication rounds. This adaptive scheme significantly enhances the robustness of our algorithm and eliminates the risk associated with selecting inappropriate penalty parameters for individual clients.
  3. 3.
    The convergence of our proposed algorithm using the inexactness criterion is analyzed. Extensive numerical experiments demonstrate the improved performance of our proposed inexactness criterion and self-adaptive penalty adjusting scheme. As validated by some numerical tests, our proposed algorithm reduces the clients’ local computational load by 64.3% while accelerating the learning process when compared to the vanilla FedADMM.
The rest of the paper is organized as follows. In Section 2, we first provide a brief background on FL and ADMM, then the vanilla application of ADMM in FL and the resulting FedADMM algorithm. Section 3 presents our proposed FedADMM-InSa method, including the inexactness criterion and the self-adaptive penalty parameter scheme. The experimental setup and simulation results are presented in Section 4. Finally, we conclude the paper in Section 5 and provide more technical details in Appendix A.
2 Preliminaries
In this section, we briefly introduce the mathematical formulation of FL and ADMM, which are central subjects of the study in this paper.
2.1 Federated Learning
Mathematically, the training process of horizontal FL can be formulated as the following minimization problem:
$\min_{z\in\mathbb{R}^{n}}\sum_{i=1}^{m}\alpha_{i}f_{i}(z),$ (2.1)
where $m$ is the number of clients, $z\in\mathbb{R}^{n}$ is the trainable parameters, $f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the local loss function of each client $i\in[m]\coloneqq\{1,\ldots,m\}$, $\alpha_{i}>0$ is the weight coefficient assigned to client $i$ by the server, and $\sum_{i=1}^{m}\alpha_{i}=1$. A common choice is $\alpha_{i}=N_{i}/N$, where $N_{i}$ is client $i$’s data volume and $N=\sum_{i=1}^{m}N_{i}$, i.e., weighting clients proportionally to their data volumes.
2.2 ADMM
Given the following constrained optimization problem:
$\min_{x\in\mathbb{R}^{n_{1}},y\in\mathbb{R}^{n_{2}}}\theta_{1}(x)+\theta_{2}(y)\text{, s.t. }Ax+By=b,$ (2.2)
where $\theta_{1}:\mathbb{R}^{n_{1}}\rightarrow\mathbb{R},\theta_{2}:\mathbb{R}^{n_{2}}\rightarrow\mathbb{R}$, $A\in\mathbb{R}^{{n_{3}}\times n_{1}},B\in\mathbb{R}^{{n_{3}}\times n_{2}}$, and $b\in\mathbb{R}^{n_{3}}$, its augmented Lagrangian function $L_{\beta}:\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}\times\mathbb{R}^{n_{3}}\rightarrow\mathbb{R}$ is defined as
$L_{\beta}(x,y,\lambda)=\theta_{1}(x)+\theta_{2}(y)-\lambda^{T}(Ax+By-b)+\frac{\beta}{2}\|Ax+By-b\|^{2},$ (2.3)
[…]
$x^{k+1}=\arg\min_{x\in\mathbb{R}^{n_{1}}}L_{\beta}\left(x,y^{k},\lambda^{k}\right),$ (2.4a)
$y^{k+1}=\arg\min_{y\in\mathbb{R}^{n_{2}}}L_{\beta}\left(x^{k+1},y,\lambda^{k}\right),$ (2.4b)
$\lambda^{k+1}=\lambda^{k}-\beta\left(Ax^{k+1}+By^{k+1}-b\right).$ (2.4c)
The ADMM is a variant of the classic augmented Lagrangian method (ALM) [17, 27], where the subproblem of each ALM iteration is decomposed into two parts and then solved in a Gauss-Seidel manner. The ADMM possesses a distinct advantage as the decomposed subproblems are typically much easier than those encountered in the ALM, allowing for the exploitation of inherent properties and structures within the model under investigation. Moreover, the ADMM usually exhibits satisfactory numerical performance without the need for specific initial iterates. These characteristics make the ADMM a benchmark algorithm across various domains, including image processing [15], statistical learning [12, 38], and optimal control problems [10, 11, 29, 39], etc. For a more comprehensive understanding of the ADMM, one can refer to [4, 8].
2.3 Vanilla FedADMM
In this subsection, we present the application of ADMM in the context of FL, leading to the derivation of the vanilla FedADMM [42]. Subsequently, we systematically analyze the challenges and limitations associated with each iterative step of the vanilla FedADMM. To apply the ADMM to address the FL problem (2.1), we first introduce auxiliary variables $u_{i}=z$, $i\in[\mathrm{m}]$, and rewrite problem (2.1) into the following consensus setting:
$\min_{u_{i},z\in\mathbb{R}^{n}}\sum_{i=1}^{m}\alpha_{i}f_{i}(u_{i}),\text{ s.t. }u_{i}=z,\forall i\in[m].$ (2.5)
It is clear that problem (2.5) is equivalent to (2.1) in the sense that their optimal solutions coincide. The formulation of problem (2.5) naturally fits into the FL setting, where $u_{i}\in\mathbb{R}^{n}$ can be interpreted as the local model parameters held by client $i\in[m]$, and $z\in\mathbb{R}^{n}$ as the global model parameters held by the server. Therefore, in the following discussions, we concentrate on solving the optimization problem (2.5) instead of (2.1), as they are equivalent. For problem (2.5), we define its augmented Lagrangian function as
$L_{\alpha,\beta}(u,\lambda,z)=\sum_{i=1}^{m}{\alpha_{i}}L_{\beta_{i}}(u_{i},\lambda_{i},z),~{}\text{with}~{}L_{\beta_{i}}(u_{i},\lambda_{i},z)=f_{i}(u_{i})-\lambda_{i}^{\top}\left(u_{i}-z\right)+\frac{\beta_{i}}{2}\left\|u_{i}-z\right\|^{2}.$ (2.6)
Here, $\lambda_{i}\in\mathbb{R}^{n}$ and $\beta_{i}>0$ are the Lagrangian multiplier and the penalty parameter for client $i\in[m]$, respectively. For simplicity, we also denote
$u=\left(u_{1},\ldots,u_{m}\right)^{\top}\in\mathbb{R}^{mn},\ \lambda=\left(\lambda_{1},\ldots,\lambda_{m}\right)^{\top}\in\mathbb{R}^{mn},$ (2.7)
$\alpha=\left(\alpha_{1},\ldots,\alpha_{m}\right)^{\top}\in\mathbb{R}^{m},\ \beta=\left(\beta_{1},\ldots,\beta_{m}\right)^{\top}\in\mathbb{R}^{m}.$
Then, given an initial point $\{z^{0},\lambda^{0}\}$, for $k\geq 0$, we can have the vanilla FedADMM that iteratively updates the variables $\left\{u,\lambda,z\right\}$ in the following steps:
$u_{i}^{k+1}$ $=\arg\min_{u_{i}\in\mathbb{R}^{n}}L_{\beta_{i}}\left(u_{i},\lambda_{i}^{k},z^{k}\right),\forall i\in[m],$ (2.8a)
$\lambda_{i}^{k+1}$ $=\lambda_{i}^{k}-\beta_{i}\left(u_{i}^{k+1}-z^{k}\right),\forall i\in[m],$ (2.8b)
$z^{k+1}$ $=\arg\min_{z\in\mathbb{R}^{n}}L_{\alpha,\beta}\left(u^{k+1},\lambda^{k+1},z\right).$ (2.8c)
As seen above, these iterative updates naturally suit the FL setting. The updates of $u_{i}^{k+1}$ in (2.8a) and $\lambda_{i}^{k+1}$ in (2.8b) represent client $i$’s local update process, and can be performed in parallel by each client $i\in[m]$. Meanwhile, the update of $z^{k+1}$ in (2.8c) is the server aggregation process after receiving $u^{k+1}$, $\lambda^{k+1}$, and $\beta$ from the clients. We then take a closer look at the solutions of the subproblems (2.8a) - (2.8c) and analyze the associated difficulties and limitations.
2.3.1 Client’s Local Update: Subproblems (2.8a) and (2.8b)
The $u$-subproblem (2.8a) and the $\lambda$-subproblem (2.8b) correspond to the client’s local update process and can be performed in parallel. The $\lambda$-subproblem (2.8b) is simple and has an analytical form. However, the $u$-subproblem (2.8a) is challenging due to the use of high-dimensional and nonlinear neural networks (such as the ResNets [16]) in FL. It generally lacks a closed-form solution. As a result, the subproblem (2.8a) should be solved iteratively and inexactly, and the implementation of the FedADMM (2.3) must be embedded by an internal iterative process for the subproblem (2.8a) In the vanilla FedADMM method, the clients normally use gradient-based methods for solving (2.8a) to a certain precision. For example, client $i$ can implement a prescribed number of $E$ steps of the full-batch gradient descent as follows:
$u_{i}^{k,e+1}=u_{i}^{k,e}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}\left(u_{i}^{k,e},\lambda_{i}^{k},z^{k}\right),\quad e=0,1,\ldots,E-1,$ (2.9)
where $\eta_{i}>0$ is the learning rate (also known as the step size). Then, client $i$ sets $u_{i}^{k+1}\coloneqq u_{i}^{k,E}$ and uses it to update $\lambda_{i}^{k+1}$ by (2.8b). While other optimization schemes like stochastic gradient descent and quasi-Newton methods such as L-BFGS [25] are plausible, the accuracy of this inexact solution significantly influences the algorithm’s effectiveness [11]. A notable mathematical problem arises concerning the determination of an appropriate inexactness criterion for solving the subproblem (2.8a) inexactly. Current FedADMM methods often require an empirically perceived precision in advance. For instance, clients may be instructed to perform a fixed number of (stochastic) gradient descent steps [1, 33] or solve the subproblem up to a constant predefined accuracy [13, 41, 42]. However, empirically assigning the same amount of local training workload is not a good strategy for heterogeneous clients with non-IID data and varying computational resources. For instance, assigning too much workload may inefficiently utilize local computational resources and cause a severe straggler effect, especially among clients with limited computational resources. Moreover, there is no necessity to pursue excessively accurate solutions for (2.8a), particularly when the iterates are still far from the solution point. Additionally, current FedADMM methods employ a fixed penalty parameter $\beta$ throughout the training process, without considering the heterogeneity of clients in FL. If the penalty parameter is inappropriately chosen at the beginning, the method’s efficiency and performance would deteriorate without remedy. To tackle the above-mentioned issues, we design a readily implementable and appropriately accurate inexactness criterion for solving subproblem (2.8a) and a self-adaptive scheme to adjust each client’s penalty parameter $\beta_{i}$ in Section 3. As a result, an inexact and self-adaptive FedADMM is proposed, enabling clients to dynamically adjust local training precision each round, eliminating the need to preset accuracy. This flexibility enhances the adaptability of our algorithm to diverse clients and datasets, saves local computational resources, and mitigates the straggler effect. In addition, the self-adaptive penalty parameter scheme improves the robustness of our algorithm by eliminating the risk associated with inappropriate pre-selection of clients’ penalty parameters.
2.3.2 Server’s Aggregation: Subproblem (2.8c)
The $z$-subproblem (2.8c) is addressed by the server to update the global parameter $z^{k+1}$. It follows from (2.6) that
$L_{\alpha,\beta}\left(u^{k+1},\lambda^{k+1},z\right)=\sum_{i=1}^{m}{\alpha_{i}}\left(f_{i}(u_{i}^{k+1})-(\lambda_{i}^{k+1})^{\top}(u_{i}^{k+1}-z)+\frac{\beta_{i}}{2}\|u_{i}^{k+1}-z\|^{2}\right),$ (2.10)
which implies that
$z^{k+1}=\arg\min_{z\in\mathbb{R}^{n}}\sum_{i=1}^{m}{\alpha_{i}}\left((\lambda_{i}^{k+1})^{\top}z+\frac{\beta_{i}}{2}\|u_{i}^{k+1}-z\|^{2}\right).$ (2.11)
Hence, $z^{k+1}$ has an analytical form given by
$z^{k+1}=\frac{1}{\sum_{i=1}^{m}{\alpha_{i}}\beta_{i}}\sum_{i=1}^{m}{\alpha_{i}}\left(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1}\right).$ (2.12)
The server can also incorporate the partial client participation strategy by selecting a subset of clients $\mathcal{M}^{k}\subseteq[m]$ to calculate updated parameters $u_{i}^{k+1}$ and $\lambda_{i}^{k+1}$ using (2.8a) and (2.8b) at the $(k+1)$-th communication round. Note from (2.12) that the clients can provide the server with calculated $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ instead of $u_{i}^{k+1}$ and $\lambda_{i}^{k+1}$, thereby reducing the communication cost. Meanwhile, for the unselected clients $i\notin\mathcal{M}^{k}$, the server can simply use their previously communicated local model parameters by setting $u_{i}^{k+1}\coloneqq u_{i}^{k}$ and $\lambda_{i}^{k+1}\coloneqq\lambda_{i}^{k}$. Finally, after receiving updated parameters $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and $\beta_{i}$ from the clients, the server can get the updated global model parameters $z^{k+1}$ by (2.12). With the above analysis of the client update and the server aggregation processes, we summarize the vanilla FedADMM algorithm based on (2.3) in Algorithm 1.
Algorithm 1 Vanilla FedADMM.
1:  Inputs: Initialize $z^{0},\lambda_{i}^{0},u_{i}^{0},\beta_{i},\alpha_{i},\eta_{i},i\in[m]$.
2:  for each communication round $k=0,1,\ldots,K-1$ do
3:     Server side: Select a subset $\mathcal{M}^{k}\subseteq[m]$ of clients and send them $z^{k}$.
4:     Client side:
5:     for each client $i\in\mathcal{M}^{k}$ in parallel do
6:        Set $u_{i}^{k,0}\coloneqq z^{k}$.
7:        for each local update step $e=0,1,\ldots,E-1$ do
8:           $u_{i}^{k,e+1}=u_{i}^{k,e}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(u_{i}^{k,e},\lambda_{i}^{k},z^{k})$.
9:        end for
10:        Set $u_{i}^{k+1}\coloneqq u_{i}^{k,E}$.
11:        Update $\lambda_{i}^{k+1}=\lambda_{i}^{k}-\beta_{i}(u_{i}^{k+1}-z^{k})$.
12:        Send $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and ${\beta_{i}}$ to the server.
13:     end for
14:     for each client $i\notin\mathcal{M}^{k}$ in parallel do
15:        Set $u_{i}^{k+1}\coloneqq u_{i}^{k}$, $\lambda_{i}^{k+1}\coloneqq\lambda_{i}^{k}$.
16:        Send $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and ${\beta_{i}}$ to the server.
17:     end for
18:     Server side:
19:     Update $z^{k+1}=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}}\sum_{i=1}^{m}\alpha_{i}(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$.
20:  end for
3 An Inexact and Self-Adaptive FedADMM
In this section, we present the design of our inexact and self-adaptive FedADMM algorithm. We first present a refined server update approach with improved stability. Then, we design an inexactness criterion for solving (2.8a) and a self-adaptive penalty parameter scheme to update $\beta$. We further show that the inexactness criterion and the self-adaptive scheme can be combined with each other.
3.1 Server’s Aggregation with Memory
From (2.12), it is evident that in the vanilla FedADMM, the update of the server’s global model parameters $z^{k+1}$ depends solely on the local parameters updated by clients in the $(k+1)$-th communication round. Due to the uncertainty arising from heterogeneous local updates, the global model parameters may undergo significant fluctuations. To address this challenge, we draw inspiration from exponential moving average methods used in several fields (see e.g., [2, 5, 6]) and propose an improved method for updating global model parameters $z^{k+1}$ as follows:
$\hat{z}^{k+1}$ $=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}}\sum_{i=1}^{m}\alpha_{i}\left(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1}\right),$ (3.1a)
$z^{k+1}$ $=\frac{1}{1+\delta}\hat{z}^{k+1}+\frac{\delta}{1+\delta}z^{k},$ (3.1b)
where $\delta>0$ controls the trade-off between stability and responsiveness in the server update process. The proposed server update strategy (3.1) improves the stability of aggregation by balancing the current and previous global model parameters. This ensures the robustness of the global model in the face of variable clients’ local updates. On the one hand, it allows for the memorization of $z^{k}$, thereby minimizing the sensitivity to outliers and leading to a more stable global model. On the other hand, the strategy maintains adaptability to the changing $\hat{z}^{k+1}$, ensuring that the global model responds appropriately to evolving local updates. Meanwhile, the proposed server update strategy (3.1) can be interpreted by augmenting the vanilla $z$-subproblem (2.8c) with an additional proximal term, given by
$z^{k+1}=\arg\min_{z\in\mathbb{R}^{n}}\ L_{\alpha,\beta}\left(u^{k+1},\lambda^{k+1},z\right)+\frac{\delta}{2}\sum_{i=1}^{m}\alpha_{i}\beta_{i}\|z-z^{k}\|^{2}.$ (3.2)
3.2 An Inexact Version of FedADMM
In this subsection, we propose an inexactness criterion for solving (2.8a) to address the challenges of presetting the client’s local update precision.
Inexactness criterion of (2.8a).
We first start with analyzing the optimality condition of the $u$-subproblem (2.8a). From (2.6), we have
$L_{\beta_{i}}\left(u_{i},\lambda_{i}^{k},z^{k}\right)=f_{i}(u_{i})-(\lambda_{i}^{k})^{\top}(u_{i}-z^{k})+\frac{\beta_{i}}{2}\|u_{i}-z^{k}\|^{2},$ (3.3)
and accordingly
$\nabla_{u_{i}}L_{\beta_{i}}\left(u_{i},\lambda_{i}^{k},z^{k}\right)=\nabla f_{i}(u_{i})-\lambda_{i}^{k}+\beta_{i}\left(u_{i}-z^{k}\right).$ (3.4)
Let $u_{i}^{k+1}$ be a solution of the subproblem (2.8a), then, it satisfies
$\nabla f_{i}(u_{i}^{k+1})-\lambda_{i}^{k}+\beta_{i}\left(u_{i}^{k+1}-z^{k}\right)=0.$ (3.5)
Based on the optimality condition (3.5), for each client $i\in[m]$, we define $e_{i}^{k}(u_{i})$ as
$e_{i}^{k}(u_{i})=\nabla f_{i}(u_{i})-\lambda_{i}^{k}+\beta_{i}\left(u_{i}-z^{k}\right).$ (3.6)
It is clear that a solution $u_{i}^{k+1}$ of the $u$-subproblem (2.8a) at the $(k+1)$-th iteration satisfies $e_{i}^{k}(u_{i}^{k+1})=0$. Hence, we can use $e_{i}^{k}(u_{i})$ as the residual for the $u$-subproblem. With the help of $e_{i}^{k}(u_{i})$, we propose the following inexactness criterion. For each client $i\in[m]$, at the $(k+1)$-th iteration, it computes $u_{i}^{k+1}$ such that
$\left\|e_{i}^{k}(u_{i}^{k+1})\right\|\leq\sigma_{i}\left\|e_{i}^{k}(u_{i}^{k})\right\|,$ (3.7)
where $\sigma_{i}$ is a given constant satisfying
$0<\sigma_{i}<\frac{\sqrt{2c_{i}}}{\sqrt{2c_{i}}+\sqrt{\beta_{i}}}<1,$ (3.8)
and $c_{i}>0$ is a parameter associated with the strong convexity constant of client $i$’s loss function $f_{i}$, i.e., $(\nabla f_{i}(x)-\nabla f_{i}(y))^{\top}(x-y)\geq c_{i}\|x-y\|^{2},\forall x,y\in\mathbb{R}^{n}$. Equivalently, the condition (3.8) can be written as
$0<\sigma_{i}<\frac{\sqrt{2}}{\sqrt{2}+\sqrt{\tilde{\beta}_{i}}}<1,$ (3.9)
where $\tilde{\beta}_{i}={\beta_{i}}/{c_{i}}$. It is evident that a smaller $\tilde{\beta}_{i}$ leads to a higher permissible value of $\sigma_{i}$, indicating that the subproblem can be solved with less precision. In practical scenarios, each client can (locally) convexify its local loss function by incorporating a regularization term. Our inexactness criterion can also be applied to the non-convex problems by using a $\tilde{\beta_{i}}>0$ in (3.9). Subsequent numerical experiments in Section 4, covering both strongly convex and non-convex problems, also underscore the feasibility of our inexactness criterion (3.7). It is noteworthy that our inexactness criterion can be assessed by each client autonomously based on its present model and penalty parameter, and it can be seamlessly executed during iterations. There is no requirement to predefine any empirically perceived constant accuracy. This gives each client the flexibility to solve its subproblem inexactly, aligning with its distinct non-IID data and computational resources. Consequently, this approach effectively mitigates the potential risk of the straggler effect that may arise with resource-constrained clients. Overall, these attributes make the inexactness criterion (3.7) straightforward to implement and more likely to result in local computational savings.
FedADMM-In: An inexact FedADMM algorithm.
Based on the discussions above, the iterative three steps of FedADMM with the inexactness criterion (3.7) are given by
$u_{i}^{k+1}$ $\approx\arg\min_{u_{i}\in\mathbb{R}^{n}}L_{\beta_{i}}\left(u_{i},\lambda_{i}^{k},z^{k}\right),\forall i\in[m],$ (3.10a)
$\lambda_{i}^{k+1}$ $=\lambda_{i}^{k}-\beta_{i}\left(u_{i}^{k+1}-z^{k}\right),\forall i\in[m],$ (3.10b)
$z^{k+1}$ $=\arg\min_{z\in\mathbb{R}^{n}}\ L_{\alpha,\beta}\left(u^{k+1},\lambda^{k+1},z\right)+\frac{\delta}{2}\sum_{i=1}^{m}\alpha_{i}\beta_{i}\|z-z^{k}\|^{2}.$ (3.10c)
We denote by FedADMM-In the FedADMM algorithm with the inexactness criterion (3.7) and summarize it in Algorithm 2.
Algorithm 2 FedADMM-In: FedADMM with the inexactness criterion (3.7).
1:  Inputs: Initialize $z^{0},\lambda_{i}^{0},u_{i}^{0},\beta_{i},\sigma_{i},\alpha_{i},\eta_{i},i\in[m]$.
2:  for each communication round $k=0,1,\ldots,K-1$ do
3:     Server side: Select a subset $\mathcal{M}^{k}\subseteq[m]$ of clients and send them $z^{k}$.
4:     Client side:
5:     for each client $i\in\mathcal{M}^{k}$ in parallel do
6:        Set $\hat{u}_{i}^{k}\coloneqq z^{k}$.
7:        while $\|e_{i}^{k}(\hat{u}_{i}^{k})\|>\sigma_{i}\|e_{i}^{k}(u_{i}^{k})\|$ do
8:           Update $\hat{u}_{i}^{k}\coloneqq\hat{u}_{i}^{k}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(\hat{u}_{i}^{k},\lambda_{i}^{k},z^{k})$.
9:        end while
10:        Set $u_{i}^{k+1}\coloneqq\hat{u}_{i}^{k}$.
11:        Update $\lambda_{i}^{k+1}=\lambda_{i}^{k}-\beta_{i}(u_{i}^{k+1}-z^{k})$.
12:        Send $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and ${\beta_{i}}$ to the server.
13:     end for
14:     for each client $i\notin\mathcal{M}^{k}$ in parallel do
15:        Set $u_{i}^{k+1}\coloneqq u_{i}^{k}$, $\lambda_{i}^{k+1}\coloneqq\lambda_{i}^{k}$.
16:        Send $(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and ${\beta_{i}}$ to the server.
17:     end for
18:     Server side:
19:     Obtain $\hat{z}^{k+1}=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}}\sum_{i=1}^{m}\alpha_{i}(\beta_{i}u_{i}^{k+1}-\lambda_{i}^{k+1})$.
20:     Update $z^{k+1}=\frac{1}{1+\delta}\hat{z}^{k+1}+\frac{\delta}{1+\delta}z^{k}$.
21:  end for
Convergence analysis.
Let us analyze the convergence of our proposed FedADMM-In algorithm under the following assumptions:
Assumption 3.1.
For all $i\in[m]$, the gradient of $f_{i}$ is $s_{i}$-Lipschitz, that is,
$\left\|\nabla f_{i}(x)-\nabla f_{i}(y)\right\|\leq s_{i}\|x-y\|,s_{i}>0,\forall x,y\in\mathbb{R}^{n}.$ (3.11)
Assumption 3.2.
For all $i\in[m]$, $f_{i}$ is $c_{i}$-strongly convex, that is,
$(\nabla f_{i}(x)-\nabla f_{i}(y))^{\top}(x-y)\geq c_{i}\|x-y\|^{2},c_{i}>0,\forall x,y\in\mathbb{R}^{n}.$ (3.12)
From the strong convexity assumption, it is easy to deduce that problem (2.1) has a unique solution, denoted by $z^{*}$. As a consequence, problem (2.5), which is equivalent to (2.1), also has a unique solution $(u^{*},z^{*})$, where $u_{i}^{*}=z^{*},\forall i\in[m]$. On the other hand, from the above assumptions, the dual problem of problem (2.5) has a unique solution $\lambda^{*}$ and the strong duality holds. In the following Theorem 3.3, we show that the sequence generated by iterative scheme (3.2) (i.e., Algorithm 2 with full client participation) converges to $(u^{*},\lambda^{*},z^{*})$.
Theorem 3.3.
Let $\{w^{k}\}=\{(u^{k},\lambda^{k},z^{k})^{\top}\}$ be the sequence generated by iterative scheme (3.2) (i.e., Algorithm 2 with full client participation). Then, we have the following assertions:
$\|e_{i}^{k}(u_{i}^{k+1})\|\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0,\quad u_{i}^{k}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}u_{i}^{*},\quad\lambda_{i}^{k}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}\lambda_{i}^{*},\quad z^{k}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}z^{*},\quad\forall i\in[m].$ (3.13)
The detailed proof is presented in Appendix A.1, which is inspired by [11, Sec. 3.3]. Although the convergence proof is conducted under the strong convexity assumption, our algorithm with the proposed inexactness criterion can also be applied to the non-convex cases by using a $\tilde{\beta_{i}}>0$ in (3.9). Moreover, the effectiveness of the algorithm with partial client participation in both strongly convex and non-convex cases is validated by the numerical experiments in Section 4.
Remark 3.4.
In practical applications, one can also use $z^{k}$ instead of $u_{i}^{k}$ on the right-hand side of the inexactness criterion (3.7). We empirically demonstrate in Section 4 that such a substitution improves the FL training performance in terms of improved model accuracy and reduced computational load. This substitution alleviates the unnecessary pursuit of the client’s local minima, which is normally inconsistent with the global minima due to the non-IID datasets. The convergence analysis of the algorithm employing this modified inexactness criterion and partial client participation serves as an interesting future study.
3.3 Self-Adaptive Penalty Parameter $\beta_{i}^{k}$
In this subsection, we present a self-adaptive scheme for adjusting the penalty parameter $\beta_{i}^{k}$ for client $i$ at communication round $k$ to further improve the performance of our FedADMM-In algorithm. This also makes the algorithm more robust to different initial choices of the penalty parameter.
Self-adaptive scheme.
As shown in (3.2), the FedADMM-In algorithm is an iterative algorithm, and an intuitive stopping criterion of the iterative updates can be $u_{i}^{k+1}-u_{i}^{k}=0$ and $\lambda_{i}^{k+1}-\lambda_{i}^{k}=0$. It follows from (3.10b) that $\lambda_{i}^{k+1}-\lambda_{i}^{k}=0$ implies that $u_{i}^{k+1}-z^{k}=0$. Inspired by the stopping criterion, we define the primal residual $p_{i}^{k}$ and the dual residual $d_{i}^{k}$ as follows:
$p_{i}^{k}$ $=\|u_{i}^{k+1}-u_{i}^{k}\|,\ i\in[m],$ (3.14a)
$d_{i}^{k}$ $=\|u_{i}^{k+1}-z^{k}\|,\ i\in[m].$ (3.14b)
To adaptively update $\beta_{i}^{k}$, we use the following scheme:
$\beta_{i}^{k+1}=\begin{cases}\beta_{i}^{k}\tau,&\text{ if }d_{i}^{k}>\mu p_{i}^{k},\\ \beta_{i}^{k}/\tau,&\text{ if }p_{i}^{k}>\mu d_{i}^{k},\\ \beta_{i}^{k},&\text{ otherwise },\end{cases}$ (3.15)
where $\mu,\tau>1$ are parameters to choose, and a typical choice can be $\mu=20$ and $\tau=2$. The essence of the self-adaptive penalty parameter scheme (3.15) lies in dynamically balancing the primal residual $p_{i}^{k}$ and the dual residual $d_{i}^{k}$. For instance, $d_{i}^{k}>\mu p_{i}^{k}$ indicates an excessive dual residual $d_{i}^{k}$. Following the definition of $d_{i}^{k}$ in (3.14b), this suggests that client $i$’s updated local model parameters $u_{i}^{k+1}$ deviate significantly from the current global model parameters $z^{k}$. Consequently, client $i$ is prompted to increase its penalty parameter $\beta_{i}^{k}$ to impose a more substantial penalty on the constraint. Conversely, a reduction in $\beta_{i}^{k}$ is recommended for client $i$ to mitigate the primal residual. Finally, the iterative steps of FedADMM with an adaptive $\beta_{i}^{k}$ are given by
$u_{i}^{k+1}$ $\approx\arg\min_{u_{i}\in\mathbb{R}^{n}}L_{\beta_{i}^{k}}\left(u_{i},\lambda_{i}^{k},z^{k}\right),\forall i\in[m],$ (3.16a)
$\lambda_{i}^{k+1}$ $=\lambda_{i}^{k}-\beta_{i}^{k}\left(u_{i}^{k+1}-z^{k}\right),\forall i\in[m],$ (3.16b)
$z^{k+1}$ $=\arg\min_{z\in\mathbb{R}^{n}}\ L_{\alpha,\beta^{k}}\left(u^{k+1},\lambda^{k+1},z\right)+\frac{\delta}{2}\sum_{i=1}^{m}\alpha_{i}\beta_{i}^{k}\|z-z^{k}\|^{2}.$ (3.16c)
Compatibility with the inexactness criterion.
We can easily extend our inexactness criterion (3.7) to make it compatible with the self-adaptive penalty parameter scheme (3.15). To accomplish this, we simply replace the fixed $\beta_{i}$ with the adaptive $\beta_{i}^{k}$ in (3.8). Then, for each client $i\in[m]$, at the $(k+1)$-th communication round, it computes $u_{i}^{k+1}$ such that
$\left\|e_{i}^{k}(u_{i}^{k+1})\right\|\leq\sigma_{i}^{k}\left\|e_{i}^{k}(u_{i}^{k})\right\|,$ (3.17)
where $\sigma_{i}^{k}$ is a given constant satisfying
$0<\sigma_{i}^{k}<\frac{\sqrt{2}}{\sqrt{2}+\sqrt{\tilde{\beta}_{i}^{k}}}<1,$ (3.18)
and $\tilde{\beta}_{i}^{k}={\beta_{i}^{k}}/{c_{i}}$ can be independently evaluated by each client $i$ based on its unique situation. Hence, the inexactness criterion (3.17) with the self-adaptive penalty parameter scheme can also be executed autonomously by each client during iterations. Moreover, this allows clients to perform personalized local update steps based on their distinct data and computational resources, effectively mitigating the risk of the straggler effect.
FedADMM-InSa: An inexact and self-adaptive FedADMM.
Based on the discussions above, we propose the FedADMM-InSa algorithm using an adaptive penalty parameter $\beta_{i}^{k}$ and present it in Algorithm 3. It brings more difficulty to prove the convergence of Algorithm 3 with a varying $\beta_{i}^{k}$ in each iteration, however, the numerical experiments in Section 4 empirically validate its robustness and improved performance.
Algorithm 3 FedADMM-InSa: FedADMM with the inexactness criterion (3.17) and the self-adaptive penalty parameter scheme (3.15).
1:  Inputs: Initialize $z^{0},\lambda_{i}^{0},u_{i}^{0},\beta_{i}^{0},\sigma_{i}^{0},c_{i},\alpha_{i},\eta_{i},i\in[m]$, $\mu>1,\tau>1$.
2:  for each communication round $k=0,1,\ldots,K-1$ do
3:     Server side: Select a subset $\mathcal{M}^{k}\subseteq[m]$ of clients and send them $z^{k}$.
4:     Client side:
5:     for each client $i\in\mathcal{M}^{k}$ in parallel do
6:        Set $\hat{u}_{i}^{k}\coloneqq z^{k}$.
7:        while $\|e_{i}^{k}(\hat{u}_{i}^{k})\|>\sigma_{i}^{k}\|e_{i}^{k}(u_{i}^{k})\|$ do
8:           Update $\hat{u}_{i}^{k}\coloneqq\hat{u}_{i}^{k}-\eta_{i}\nabla_{u_{i}}L_{\beta_{i}}(\hat{u}_{i}^{k},\lambda_{i}^{k},z^{k})$.
9:        end while
10:        Set $u_{i}^{k+1}\coloneqq\hat{u}_{i}^{k}$.
11:        Update $\lambda_{i}^{k+1}=\lambda_{i}^{k}-\beta_{i}^{k}(u_{i}^{k+1}-z^{k})$.
12:        Send $(\beta_{i}^{k}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and ${\beta_{i}^{k}}$ to the server.
13:        Calculate $p_{i}^{k}=\|u_{i}^{k+1}-u_{i}^{k}\|$ and $d_{i}^{k}=\|u_{i}^{k+1}-z^{k}\|$.
14:        if $d_{i}^{k}>\mu p_{i}^{k}$ then
15:           Update $\beta_{i}^{k+1}=\beta_{i}^{k}\tau$.
16:        else if $p_{i}^{k}>\mu d_{i}^{k}$ then
17:           Update $\beta_{i}^{k+1}=\beta_{i}^{k}/\tau$.
18:        else
19:           Update $\beta_{i}^{k+1}=\beta_{i}^{k}$.
20:        end if
21:        Update $\sigma_{i}^{k+1}=\frac{\sqrt{2}}{\sqrt{2}+\sqrt{{\beta_{i}^{k+1}}/{c_{i}}}}$.
22:     end for
23:     for each client $i\notin\mathcal{M}^{k}$ in parallel do
24:        Set $u_{i}^{k+1}\coloneqq u_{i}^{k}$, $\lambda_{i}^{k+1}\coloneqq\lambda_{i}^{k}$, $\beta_{i}^{k+1}\coloneqq\beta_{i}^{k}$, $\sigma_{i}^{k+1}\coloneqq\sigma_{i}^{k}$.
25:        Send $(\beta_{i}^{k}u_{i}^{k+1}-\lambda_{i}^{k+1})$ and ${\beta_{i}^{k}}$ to the server.
26:     end for
27:     Server side:
28:     Obtain $\hat{z}^{k+1}=\frac{1}{\sum_{i=1}^{m}\alpha_{i}\beta_{i}^{k}}\sum_{i=1}^{m}\alpha_{i}(\beta_{i}^{k}u_{i}^{k+1}-\lambda_{i}^{k+1})$.
29:     Update $z^{k+1}=\frac{1}{1+\delta}\hat{z}^{k+1}+\frac{\delta}{1+\delta}z^{k}$.
30:  end for
4 Experimental Results
In this section, we conduct extensive numerical tests to demonstrate the improved performance of our proposed FedADMM-In (Algorithm 2) and FedADMM-InSa (Algorithm 3) in comparison to the vanilla FedADMM (Algorithm 1). We first introduce the experimental settings and the implementation details used in our experiments. Then, we present the simulation results and analysis.
4.1 Setups
Example 1: Linear regression with a synthetic dataset.
In this example, we set each client’s local loss function to be
$f_{i}(u_{i})=\frac{1}{2N_{i}}\sum_{j=1}^{N_{i}}\left(u_{i}^{\top}a_{i}^{j}-b_{i}^{j}\right)^{2}+\frac{\gamma_{i}}{2}\|u_{i}\|^{2},\quad i\in[m].$ (4.1)
[…]
Example 2: Image classification with the MNIST dataset.
In the second example, we address the image classification problem using fully-connected neural networks with the MNIST dataset [20]. The MNIST dataset contains images of handwritten digits. The samples in this dataset are $28\times 28$ grayscale images with handwritten digits from 0 to 9 in the center. We select a training set of 10,000 examples and a test set of 1,000 examples. The number of clients is set to $m=100$. For the neural networks, each client uses a fully-connected neural network with two hidden layers, each comprising 200 neurons, incorporating the cross-entropy loss and the ReLU activation function.
Non-IID data separation.
To test the performance of the algorithms on non-IID data, we separate datasets in the following way. In both examples, there are in total $N=10,000$ training samples and $m=100$ clients. In Example 1, we first shuffle the training samples and then distribute them evenly among all clients. In Example 2, we first sorted the data by labels and then divided them into several shards, each shard containing images with the same label. Then, for 100 clients, the training dataset with 10,000 samples is divided into 200 shards of size 50, and each client receives 2 shards. This means each client has at most two kinds of labels in their dataset, corresponding to a pathological non-IID scenario.
FL training parameters.
In each communication round, the server uniformly samples 10% of the clients to perform the local training. During the local training, the selected clients use full-batch gradient descent with a learning rate of 0.01. For the vanilla FedADMM algorithm, unless otherwise indicated, the clients perform fixed $E=10$ steps of the full-batch gradient descent. For FedADMM-In and FedADMM-InSa algorithms, the number of the full-batch gradient descent steps is controlled by the inexactness criterion, and we constrain the maximum number of gradient descent steps in each communication round to 10. For the inexactness criterion, we set $\tilde{\beta}_{i}$ to the same value as ${\beta_{i}}$ and replace $u_{i}^{k}$ with $z^{k}$ for all the tests. We use the weight coefficient $\alpha_{i}=0.01$ because the amount of data per client is the same in both examples. We choose $\delta=0.01$ for the server aggregation. Finally, important parameters used in the tests are summarized in Table 1.
Table 1: Dataset parameters and FL training parameters.
Parameters
Example 1
Linear regression
Example 2
Image classification
Training set size 10000 10000
Test set size - 1000
Data dimension 1000 $28\times 28$
Communication rounds $K$ 100 200
Number of clients $m$ 100 100
Client subset size $|\mathcal{M}^{k}|$ 10 10
Client learning rate $\eta_{i}$ 0.01 0.01
4.2 Evaluation of FedADMM-In (Algorithm 2) with Example 1
In this subsection, we present the comparison results of our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 1. The results are plotted in Figure 1. In this experiment, we test scenarios with $\beta_{i}=1$ and $\beta_{i}=2$. In the case of $\beta_{i}=2$, as shown in Figure 0(a), FedADMM-In achieves lower loss values from the beginning and maintains this advantage throughout the experiment. For $\beta_{i}=1$, although FedADMM-In initially exhibits a slower loss decline than FedADMM, it reaches a similar loss value to FedADMM after about 30 rounds. Meanwhile, as illustrated in Figure 0(b) and Figure 0(c), FedADMM-In requires significantly fewer local gradient descent steps than FedADMM. These results indicate that our proposed FedADMM-In algorithm can achieve comparable or even superior training outcomes while reducing the computational load on clients.
(a) Training loss.
(b) Client 1’s local update steps.
(c) Client 1’s local update steps.
Figure 1: Comparison results of our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 1. For the vanilla FedADMM, the clients perform fixed $E=10$ local full-batch gradient descent steps. For our FedADMM-In, the number of full-batch gradient descent steps is controlled by the inexactness criterion (3.7).
4.3 Evaluation of FedADMM-In (Algorithm 2) with Example 2
In this subsection, we conduct the comparison tests between our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 2 and present the results in Figure 2. In both sets of experiments conducted with penalty parameters $\beta_{i}=1$ and $\beta_{i}=2$, our proposed FedADMM-In consistently demonstrates superior performance. This is evidenced by achieving higher prediction accuracy and lower loss after the same 200 communication rounds, as illustrated in Figure 1(a) and Figure 1(b), respectively. Additionally, we visualize the number of local gradient descent steps performed by a specific client (client 1) in each communication round, represented in Figure 1(c) and Figure 1(d). It is evident that employing our proposed FedADMM-In algorithm allows clients to execute fewer gradient descent steps as well as ensure training efficiency. This significantly reduces the potential waste of valuable computational resources in FL. However, it is noteworthy from Figure 2 that when $\beta_{i}=2$, the performance of both algorithms deteriorates compared to the case of $\beta_{i}=1$. This discrepancy reflects the substantial impact of the initial pre-selected penalty parameters on the training performance. To address this issue, we propose the FedADMM-InSa algorithm, the performance of which is validated in the next subsection.
(a) Test accuracy.
(b) Training loss.
(c) Client 1’s local update steps ($\beta_{i}=1$).
(d) Client 1’s local update steps ($\beta_{i}=2$).
Figure 2: Comparison results of our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 2. For the vanilla FedADMM, the clients perform fixed $E=10$ local full-batch gradient descent steps. For our FedADMM-In, the number of the full-batch gradient descent steps is controlled by the inexactness criterion (3.7).
4.4 Evaluation of FedADMM-InSa (Algorithm 3) with Example 2
In this subsection, we present the comparison results of our FedADMM-InSa (Algorithm 3) with the other two algorithms, namely our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1). For FedADMM-InSa, we set the initial penalty parameter $\beta_{i}^{0}=2$ and update it by the scheme (3.15) with $\mu=20$ and $\tau=2$. For both FedADMM-In and FedADMM, the penalty parameter $\beta_{i}=2$ is fixed throughout the training. For the vanilla FedADMM, we test it with three different local full-batch gradient descent steps, i.e., $E=2,5,10$. The results are plotted in Figure 3 and summarized in Table 2. From Figure 3, it is evident that our FedADMM-InSa algorithm, incorporating a self-adaptive penalty parameter (3.15), significantly outperforms the other two algorithms. As depicted in Figure 2(a), under the same number of communication rounds, FedADMM-InSa exhibits significantly higher prediction accuracy compared to the other two algorithms. In other words, FedADMM-InSa can achieve a given target accuracy with fewer communication rounds than the other two algorithms. This is also validated by the fastest decreasing training loss of FedADMM-InSa shown in Figure 2(b). Figure 2(d) illustrates that for a specific client (client 1), when using FedADMM-InSa, it performs less number of gradient descent steps compared with FedADMM-In and FedADMM ($E=10$), especially in the later communication rounds. This reduction in gradient descent steps is further quantified in Table 2, where compared to the baseline FedADMM ($E=10$) algorithm, our proposed FedADMM-In and FedADMM-InSa algorithms reduce the number of gradient descent steps by 49.8% and 64.3%, respectively, leading to a substantial improvement in training efficiency. Although FedADMM ($E=2$) has the lowest total number of local gradient descent steps, the value of $E$ has to be chosen empirically in advance. Without knowing the empirical information of the FL system, it is difficult to find an appropriate $E$ in advance. Meanwhile, the decline in test accuracy from FedADMM ($E=2$) to FedADMM ($E=10$) is likely attributed to the pathological non-IID datasets, further emphasizing the challenge of empirically choosing an appropriate $E$ in advance. Notably, our proposed FedADMM-InSa algorithm outperforms FedADMM ($E=2$) in terms of test accuracy, achieving a balance between improving training performance and reducing computational load. Finally, in Figure 2(c), the variation of the average penalty parameter across all clients is plotted. It can be observed that FedADMM-InSa dynamically adjusts the average penalty parameter from an initial value of two to around one. This demonstrates that FedADMM-InSa can adaptively correct the initially improperly chosen penalty parameters, highlighting the robustness of our proposed algorithm.
(a) Test accuracy.
(b) Training loss.
(c) Average of all clients’ penalty parameters.
(d) Client 1’s local update steps.
Figure 3: Comparison results of our FedADMM-InSa (Algorithm 3), our FedADMM-In (Algorithm 2) and the vanilla FedADMM (Algorithm 1) with Example 2. For our FedADMM-InSa, the initial penalty parameter $\beta_{i}^{0}=2$ and updates by scheme (3.15) with $\mu=20$ and $\tau=2$. For the vanilla FedADMM, the client’s local full-batch gradient descent steps $E=2,5,10$. For both FedADMM-In and FedADMM, the penalty parameter is fixed as $\beta_{i}=2$ throughout the training.
Table 2: Detailed comparison results of the tests presented in Figure 3. The results in the steps reduction column use the vanilla FedADMM ($E=10$) as the baseline.
Algorithm Test accuracy Training loss Local update steps Steps reduction
FedADMM ($E=2$) 81.6% 0.528 4000 -
FedADMM ($E=5$) 71.9% 0.762 10000 -
FedADMM ($E=10$) 62.6% 0.955 20000 0.0%
FedADMM-In (ours) 70.9% 0.730 10036 49.8%
FedADMM-InSa (ours) 87.8% 0.325 7139 64.3%
5 Conclusions
In this paper, we introduce the FedADMM-InSa algorithm, a novel approach that leverages the alternating direction method of multipliers (ADMM) to address the challenges of federated learning (FL) in the presence of data and system heterogeneity. Distinguished from current FedADMM methods, our inexact and self-adaptive algorithm mitigates the need for intricate empirical hyperparameter settings. The introduced inexactness criterion improves computational efficiency by removing the requirement to determine local training accuracy in advance. Additionally, the self-adaptive scheme dynamically adjusts each client’s penalty parameter, enhancing the robustness of our algorithm. Numerical tests demonstrate the reduction in clients’ local computational load and accelerated learning performance on both synthetic and real-world datasets, highlighting the practical advancements our approach brings to FL systems. In addition to the demonstrated benefits of our proposed algorithm, the integration of privacy-preserving techniques and the investigation of the algorithm’s performance in large-scale FL applications are interesting and challenging future directions.
A Appendix
A.1 Convergence Analysis of the Sequence Generated by (3.2)
A.1.1 Notations
To present our convergence analysis in a compact form, we first rewrite the augmented Lagrangian function (2.6) as
$L_{\alpha,\beta}(u,\lambda,z)$ $=\sum_{i=1}^{m}{\alpha_{i}}\left(f_{i}(u_{i})-\lambda_{i}^{\top}\left(u_{i}-z\right)+\frac{\beta_{i}}{2}\left\|u_{i}-z\right\|^{2}\right)$ (A.1)
$=\alpha^{\top}{f}(u)-\lambda^{\top}I_{\alpha}(u-Bz)+\frac{1}{2}\|u-Bz\|_{I\alpha I_{\beta}}^{2}.$
Here, $\lambda_{i}\in\mathbb{R}^{n}$ is the Lagrange multiplier associated with the equality constraint $u_{i}=z$, $\beta_{i}>0$ is a penalty parameter,
$u=\left(u_{1},\ldots,u_{m}\right)^{\top}\in\mathbb{R}^{mn},\ \lambda=\left(\lambda_{1},\ldots,\lambda_{m}\right)^{\top}\in\mathbb{R}^{mn},$ (A.2)
$\alpha=\left(\alpha_{1},\ldots,\alpha_{m}\right)^{\top}\in\mathbb{R}^{m},\ \beta=\left(\beta_{1},\ldots,\beta_{m}\right)^{\top}\in\mathbb{R}^{m},$
$f(u)\coloneqq\left(f_{1}\left(u_{1}\right),\ldots,f_{m}\left(u_{m}\right)\right)^{\top}\in\mathbb{R}^{m},$
$\nabla f(u)\coloneqq\left(\nabla f_{1}\left(u_{1}\right),\ldots,\nabla f_{m}\left(u_{m}\right)\right)^{\top}\in\mathbb{R}^{mn},$
$e^{k}(u)\coloneqq(e_{1}^{k}(u_{1}),\ldots,e_{m}^{k}(u_{m}))^{\top}\in\mathbb{R}^{mn},$
$B=(\underbrace{I_{n},\ldots,I_{n}}_{m})^{\top}\in\mathbb{R}^{mn\times n},$
and diagonal matrices $I_{\alpha},I_{\beta}\in\mathbb{R}^{mn\times mn}$ are
$I_{\alpha}=\left(\begin{array}[]{lll}\alpha_{1}I_{n}&&\\ &\ddots&\\ &&\alpha_{m}I_{n}\end{array}\right),\ I_{\beta}=\left(\begin{array}[]{lll}\beta_{1}I_{n}&&\\ &\ddots&\\ &&\beta_{m}I_{n}\end{array}\right).$ (A.3)
We define the $H$-norm with a symmetric and positive-definite matrix $H\in\mathbb{R}^{n\times n}$ as
$\|x\|_{H}=(x^{\top}Hx)^{1/2},\quad\forall x\in\mathbb{R}^{n}.$ (A.4)
We also denote $w\in W\coloneqq\mathbb{R}^{mn}\times\mathbb{R}^{mn}\times\mathbb{R}^{n},v\in V\coloneqq\mathbb{R}^{mn}\times\mathbb{R}^{n}$ and the function $F(w)$ as follows:
$\texttt{}w=\left(\begin{array}[]{c}u\\ \lambda\\ z\end{array}\right),\ v=\left(\begin{array}[]{c}\lambda\\ z\end{array}\right),\ F(w)=\left(\begin{array}[]{c}I_{\alpha}(\nabla f(u)-\lambda)\\ I_{\alpha}(u-Bz)\\ B^{\top}I_{\alpha}\lambda\end{array}\right).$ (A.5)
Recall that $(u^{*},z^{*})$ is the unique solution of problem (2.5) and $\lambda^{*}$ is the solution of its dual problem. Let $w^{*}=(u^{*},\lambda^{*},z^{*})^{\top}$, then the first-order optimality conditions of problem (2.5) and its dual problem can be expressed as:
$F(w^{*})=0.$ (A.6)
From Assumption 3.2 and the formula of $F$, it is easy to deduce that the solution $w^{*}$ is unique.
A.1.2 Optimality Conditions
Using the notations in Appendix A.1.1, we can rewrite scheme (3.2) as
$u^{k+1}\approx\arg\min_{u\in\mathbb{R}^{mn}}L_{\alpha,\beta}\left(u,\lambda^{k},z^{k}\right),$ (A.7a)
$\lambda^{k+1}=\lambda^{k}-I_{\beta}\left(u^{k+1}-Bz^{k}\right),$ (A.7b)
$z^{k+1}=\arg\min_{z\in\mathbb{R}^{n}}L_{\alpha,\beta}\left(u^{k+1},\lambda^{k+1},z\right)+\frac{\delta}{2}\left\|B(z-z^{k})\right\|_{I_{\alpha}I_{\beta}}^{2}.$ (A.7c)
For the point $w^{k+1}=\left(u^{k+1},z^{k+1},\lambda^{k+1}\right)^{\top}$ generated by (A.1.2), utilizing the definitions of $e_{i}^{k}(u_{i})$ in (3.6) and $e^{k}(u)$ in (A.2), and combining the first-order optimality condition of (A.7c), we have:
$\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},\lambda^{k},z^{k}\right)=I_{\alpha}\left(e^{k}(u^{k+1})\right),$ (A.8a)
$\lambda^{k+1}=\lambda^{k}-I_{\beta}\left(u^{k+1}-Bz^{k}\right),$ (A.8b)
$\left(z-z^{k+1}\right)^{\top}\left(B^{\top}I_{\alpha}\lambda^{k+1}-B^{\top}I_{\alpha}I_{\beta}(u^{k+1}-Bz^{k+1})+\delta B^{\top}I_{\alpha}I_{\beta}B(z^{k+1}-z^{k})\right)\geq 0.$ (A.8c)
Moreover, note that the inexactness condition (3.8) implies that
$0<\frac{\sigma_{i}^{2}\beta_{i}}{2c_{i}(1-\sigma_{i})^{2}}=\left(\frac{\sigma_{i}}{2c_{i}(1-\sigma_{i})}\right)\left(\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}\right)<1,\quad\forall i\in[m].$ (A.9)
Then, there exists a constant $\mu_{i}>0$ such that
$\left(c_{i}-\frac{\mu_{i}}{2}\frac{\sigma_{i}}{1-\sigma_{i}}\right)>0\quad\text{ and }\quad\left(1-\frac{1}{\mu_{i}}\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}\right)>0,\quad\forall i\in[m].$ (A.10)
These inequalities will be used later in the proof.
A.1.3 Convergence Proof
With the above preparations, we start to prove the convergence of sequence $\left\{w^{k}\right\}$ generated by (3.2). We first prove two lemmas which will be useful in the following discussion. First of all, we analyze how different the point $w^{k}$ generated by our algorithm is away from the solution $w^{*}$ of (A.6).
Lemma A.1.
Let $\{w^{k}\}=\{(u^{k},\lambda^{k},z^{k})^{\top}\}$ be the sequence generated by scheme (3.2). Then, for all $w\in W$, one has
$\left(w^{k+1}-w\right)^{\top}F\left(w^{k+1}\right)\leq$ $\left(u^{k+1}-u\right)^{\top}\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},\lambda^{k},z^{k}\right)$ (A.11)
$+\frac{1}{2}\left(\|v^{k}-v\|_{H_{1}}^{2}-\|v^{k+1}-v\|_{H_{1}}^{2}-\|v^{k}-v^{k+1}\|_{H_{1}}^{2}\right),$
where
$v=\left(\begin{array}[]{l}\lambda\\ z\end{array}\right),\ H_{1}=\left(\begin{array}[]{ll}I_{\alpha}I_{\beta}^{-1}&I_{\alpha}B\\ B^{\top}I_{\alpha}&(1+\delta)B^{\top}I_{\alpha}I_{\beta}B\end{array}\right)\succ 0.$ (A.12)
Proof.
For all $w\in W$, we have
[…]
$+\left(\lambda^{k+1}-\lambda\right)^{\top}I_{\alpha}\left(u^{k+1}-Bz^{k+1}\right)$
$+\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}\lambda^{k+1}$
It follows from (A.1) that
$\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},\lambda^{k},z^{k}\right)$ $=I_{\alpha}\left(\nabla f(u^{k+1})-\lambda^{k}+I_{\beta}(u^{k+1}-Bz^{k})\right)$ (A.14)
=(A.8b)${=}I_{\alpha}\left(\nabla f(u^{k+1})-\lambda^{k+1}\right).$ ) .
For the first term on the right-hand side of (A.13), it follows from (A.14) that
[…]
For the second term on the right-hand side of (A.13), we have
$\left(\lambda^{k+1}-\lambda\right)^{\top}I_{\alpha}\left(u^{k+1}-Bz^{k+1}\right)$ (A.16)
$=$ $\left(\lambda^{k+1}-\lambda\right)^{\top}I_{\alpha}\left(u^{k+1}-Bz^{k}+Bz^{k}-Bz^{k+1}\right)$
$=$ $\left(\lambda^{k+1}-\lambda\right)^{\top}I_{\alpha}\left(u^{k+1}-Bz^{k}\right)+\left(\lambda^{k+1}-\lambda\right)^{\top}I_{\alpha}B\left(z^{k}-z^{k+1}\right)$
${=}$ $\left(\lambda^{k+1}-\lambda\right)^{\top}I_{\alpha}I_{\beta}^{-1}\left(\lambda^{k}-\lambda^{k+1}\right)+\left(\lambda^{k+1}-\lambda\right)^{\top}I_{\alpha}B\left(z^{k}-z^{k+1}\right)$
For the third term on the right-hand side of (A.13), we have
$\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}\lambda^{k+1}$ (A.17)
(${\leq}$ $\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}I_{\beta}\left(u^{k+1}-Bz^{k+1}\right)+\delta\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}I_{\beta}B\left(z^{k}-z^{k+1}\right)$
$=$ $\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}I_{\beta}\left(u^{k+1}-Bz^{k}+Bz^{k}-Bz^{k+1}\right)$
$+\left(z^{k+1}-z\right)^{\top}\delta B^{\top}I_{\alpha}I_{\beta}B\left(z^{k}-z^{k+1}\right)$
$=$ $\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}I_{\beta}\left(u^{k+1}-Bz^{k}\right)+\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}I_{\beta}B\left(z^{k}-z^{k+1}\right)$
$+\delta\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}I_{\beta}B\left(z^{k}-z^{k+1}\right)$
${=}$ $\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}\left(\lambda^{k}-\lambda^{k+1}\right)+\left(1+\delta\right)\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}I_{\beta}B\left(z^{k}-z^{k+1}\right).$
Substituting (A.15)-(A.17) back into (A.13) yields
$\left(w^{k+1}-w\right)^{\top}F\left(w^{k+1}\right)$ (A.18)
$\leq$ $\left(u^{k+1}-u\right)^{\top}\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},\lambda^{k},z^{k}\right)$
$+\left(\lambda^{k+1}-\lambda\right)^{\top}I_{\alpha}I_{\beta}^{-1}\left(\lambda^{k}-\lambda^{k+1}\right)+\left(\lambda^{k+1}-\lambda\right)^{\top}I_{\alpha}B\left(z^{k}-z^{k+1}\right)$
$+\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}\left(\lambda^{k}-\lambda^{k+1}\right)+\left(1+\delta\right)\left(z^{k+1}-z\right)^{\top}B^{\top}I_{\alpha}I_{\beta}B\left(z^{k}-z^{k+1}\right).$
Using $v$ and $H_{1}$ defined by (A.12), we can rewrite (A.18) in a more compact form:
$\left(w^{k+1}-w\right)^{\top}F\left(w^{k+1}\right)\leq$ $\left(u^{k+1}-u\right)^{\top}\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},\lambda^{k},z^{k}\right)$ (A.19)
$+\left(v^{k+1}-v\right)^{\top}H_{1}\left(v^{k}-v^{k+1}\right).$
Applying the identity
$(c-a)^{\top}M(b-c)=\frac{1}{2}\left(\|b-a\|_{M}^{2}-\|c-a\|_{M}^{2}-\|b-c\|_{M}^{2}\right)$ (A.20)
to the second term on the right-hand side of (A.19), we have
$\left(v^{k+1}-v\right)^{\top}H_{1}\left(v^{k}-v^{k+1}\right)=\frac{1}{2}\left(\|v^{k}-v\|_{H_{1}}^{2}-\|v^{k+1}-v\|_{H_{1}}^{2}-\|v^{k}-v^{k+1}\|_{H_{1}}^{2}\right).$ (A.21)
Substituting (A.21) back into (A.19) yields
$\left(w^{k+1}-w\right)^{\top}F\left(w^{k+1}\right)\leq$ $\left(u^{k+1}-u\right)^{\top}\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},\lambda^{k},z^{k}\right)$ (A.22)
$+\frac{1}{2}\left(\|v^{k}-v\|_{H_{1}}^{2}-\|v^{k+1}-v\|_{H_{1}}^{2}-\|v^{k}-v^{k+1}\|_{H_{1}}^{2}\right).$
We thus complete the proof. ∎
On the right-hand side of (A.11), the three quadratic terms are easily amenable to manipulation across various indicators through algebraic operations. However, it is less apparent how the cross-term can be controlled to demonstrate the convergence of the sequence $\{w^{k}\}$. Therefore, we study this term and demonstrate that the sum of these cross-terms over $K\geq 1$ iterations can be bounded by certain quadratic terms. This result is shown in the following Lemma A.2.
Lemma A.2.
Let $\{w^{k}\}=\{(u^{k},\lambda^{k},z^{k})^{\top}\}$ be the sequence generated by scheme (3.2). For any integer $K\geq 1$ and $\mu_{i}$ satisfying (A.10), one has
$\sum_{k=1}^{K}\left(u^{k+1}-u\right)^{\top}\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},\lambda^{k},z^{k}\right)$ (A.23)
$\leq$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\frac{\mu_{i}}{2}\frac{\sigma_{i}}{1-\sigma_{i}}\left\|u_{i}^{k+1}-u_{i}\right\|^{2}+\sum_{k=1}^{K-1}\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}}{1-\sigma_{i}}\beta_{i}\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{H_{\beta_{i}}}^{2}$
$+\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2},\quad\forall u\in\mathbb{R}^{mn},$
where
$v_{i}=\left(\begin{array}[]{l}\lambda_{i}\\ z\end{array}\right),\ H_{\beta_{i}}=\left(\begin{array}[]{cc}\frac{1}{\beta_{i}}I_{n}&I_{n}\\ I_{n}&\left(1+\delta\right)\beta_{i}I_{n}\end{array}\right)\succ 0.$ (A.24)
Proof.
For the residual $e_{i}^{k}(u_{i}),i\in[m]$, it follows from (3.6) that
$e_{i}^{k}\left(u_{i}^{k}\right)$ $=\nabla f_{i}\left(u_{i}^{k}\right)-\lambda_{i}^{k}+\beta_{i}\left(u_{i}^{k}-z^{k}\right),$ (A.25a)
$e_{i}^{k-1}\left(u_{i}^{k}\right)$ $=\nabla f_{i}\left(u_{i}^{k}\right)-\lambda_{i}^{k-1}+\beta_{i}\left(u_{i}^{k}-z^{k-1}\right).$ (A.25b)
Combining (A.25a) and (A.25b) we have
$e_{i}^{k}(u_{i}^{k})=e_{i}^{k-1}(u_{i}^{k})+\beta_{i}(z^{k-1}-z^{k})+\lambda_{i}^{k-1}-\lambda_{i}^{k}.$ (A.26)
For the residual $e_{i}^{k}(u_{i}^{k+1}),i\in[m]$, it follows from (3.7) that
[…]
${=}$ $\sigma_{i}\left\|e_{i}^{k-1}(u_{i}^{k})+\beta_{i}(z^{k-1}-z^{k})+\lambda_{i}^{k-1}-\lambda_{i}^{k}\right\|$
$\leq$ $\sigma_{i}\left\|e_{i}^{k-1}(u_{i}^{k})\right\|+\sigma_{i}\left\|\beta_{i}(z^{k}-z^{k-1})+\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|.$
For the second term on the right-hand side of (A.27), with any $\delta\geq 0$, we have
$\sigma_{i}\left\|\beta_{i}(z^{k}-z^{k-1})+\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|$ (A.28)
$=$ $\sigma_{i}\left(\beta_{i}^{2}\left\|z^{k}-z^{k-1}\right\|^{2}+\left\|\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|^{2}+2\beta_{i}\left(z^{k}-z^{k-1}\right)^{\top}\left(\lambda_{i}^{k}-\lambda_{i}^{k-1}\right)\right)^{\frac{1}{2}}$
$=$ $\sigma_{i}\sqrt{\beta_{i}}\left(\beta_{i}\left\|z^{k}-z^{k-1}\right\|^{2}+\frac{1}{\beta_{i}}\left\|\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|^{2}+2\left(z^{k}-z^{k-1}\right)^{\top}\left(\lambda_{i}^{k}-\lambda_{i}^{k-1}\right)\right)^{\frac{1}{2}}$
$\leq$ $\sigma_{i}\sqrt{\beta_{i}}\left(\left(1+\delta\right)\beta_{i}\left\|z^{k}-z^{k-1}\right\|^{2}+\frac{1}{\beta_{i}}\left\|\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|^{2}+2\left(z^{k}-z^{k-1}\right)^{\top}\left(\lambda_{i}^{k}-\lambda_{i}^{k-1}\right)\right)^{\frac{1}{2}}.$
Using the $H$-norm notation in (A.4) with $v_{i}$ and $H_{\beta_{i}}$ defined by (A.24), we have
$\left\|v_{i}^{k}-v_{i}^{k-1}\right\|_{H_{\beta_{i}}}^{2}=\left(1+\delta\right)\beta_{i}\left\|z^{k}-z^{k-1}\right\|^{2}+\frac{1}{\beta_{i}}\left\|\lambda_{i}^{k}-\lambda_{i}^{k-1}\right\|^{2}+2\left(z^{k}-z^{k-1}\right)^{\top}\left(\lambda_{i}^{k}-\lambda_{i}^{k-1}\right).$ (A.29)
Substituting (A.28) back into (A.27) and using (A.29), we have
$\left\|e_{i}^{k}(u_{i}^{k+1})\right\|\leq$ $\sigma_{i}\left\|e_{i}^{k-1}(u_{i}^{k})\right\|+\sigma_{i}\sqrt{\beta_{i}}\left\|v_{i}^{k}-v_{i}^{k-1}\right\|_{H_{\beta_{i}}}$ (A.30)
$\leq$ $\sum_{j=0}^{k-1}\sigma_{i}^{k-j}\sqrt{\beta_{i}}\left\|v_{i}^{j}-v_{i}^{j+1}\right\|_{H_{\beta_{i}}}+\sigma_{i}^{k}\left\|e_{i}^{0}(u_{i}^{1})\right\|.$
By using (A.8a) and (A.30), for any $\mu_{i}>0$ satisfying (A.10) and $u\in\mathbb{R}^{mn}$, we have
[…]
$=$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\left(u_{i}^{k+1}-u_{i}\right)^{\top}e_{i}^{k}(u_{i}^{k+1})\leq\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\left\|u_{i}^{k+1}-u_{i}\right\|\left\|e_{i}^{k}(u_{i}^{k+1})\right\|$
(${\leq}$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\sum_{j=0}^{k-1}\alpha_{i}\sigma_{i}^{k-j}\sqrt{\beta_{i}}\left\|u_{i}^{k+1}-u_{i}\right\|\left\|v_{i}^{j}-v_{i}^{j+1}\right\|_{H_{\beta_{i}}}+\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\sigma_{i}^{k}\left\|u_{i}^{k+1}-u_{i}\right\|\left\|e_{i}^{0}(u_{i}^{1})\right\|$
$\leq$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\sum_{j=1}^{k-1}\alpha_{i}\sigma_{i}^{k-j}\sqrt{\beta_{i}}\left\|u_{i}^{k+1}-u_{i}\right\|\left\|v_{i}^{j}-v_{i}^{j+1}\right\|_{H_{\beta_{i}}}$
$+\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\sigma_{i}^{k}\left\|u_{i}^{k+1}-u_{i}\right\|\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)$ (A.31)
$\leq$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\sum_{j=1}^{k-1}\frac{\mu_{i}}{2}\alpha_{i}\sigma_{i}^{k-j}\left\|u_{i}^{k+1}-u_{i}\right\|^{2}+\sum_{k=1}^{K}\sum_{i=1}^{m}\sum_{j=1}^{k-1}\frac{1}{2\mu_{i}}\alpha_{i}\sigma_{i}^{k-j}\beta_{i}\left\|v_{i}^{j}-v_{i}^{j+1}\right\|_{H_{\beta_{i}}}^{2}$
$+\sum_{k=1}^{K}\sum_{i=1}^{m}\frac{\mu_{i}}{2}\alpha_{i}\sigma_{i}^{k}\left\|u_{i}^{k+1}-u_{i}\right\|^{2}+\sum_{k=1}^{K}\sum_{i=1}^{m}\frac{1}{2\mu_{i}}\alpha_{i}\sigma_{i}^{k}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}$
$=$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\sum_{j=0}^{k-1}\frac{\mu_{i}}{2}\alpha_{i}\sigma_{i}^{k-j}\left\|u_{i}^{k+1}-u_{i}\right\|^{2}+\sum_{k=1}^{K}\sum_{i=1}^{m}\sum_{j=1}^{k-1}\frac{1}{2\mu_{i}}\alpha_{i}\sigma_{i}^{k-j}\beta_{i}\left\|v_{i}^{j}-v_{i}^{j+1}\right\|_{H_{\beta_{i}}}^{2}$
$+\sum_{k=1}^{K}\sum_{i=1}^{m}\frac{1}{2\mu_{i}}\alpha_{i}\sigma_{i}^{k}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}.$
It follows from (3.8) that $\sigma_{i}\in(0,1),\forall i\in[m]$. Then, by using the property of the geometric series, we can have
$\sum_{k=1}^{K}\left(u^{k+1}-u\right)^{\top}\nabla_{u}L_{\alpha,\beta}\left(u^{k+1},z^{k},\lambda^{k}\right)$ (A.32)
$\leq$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\frac{\mu_{i}}{2}\frac{\sigma_{i}-\sigma_{i}^{k+1}}{1-\sigma_{i}}\left\|u_{i}^{k+1}-u_{i}\right\|^{2}+\sum_{i=1}^{m}\sum_{j=1}^{K-1}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}-\sigma_{i}^{K-j+1}}{1-\sigma_{i}}\beta_{i}\left\|v_{i}^{j}-v_{i}^{j+1}\right\|_{H_{\beta_{i}}}^{2}$
$+\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}-\sigma_{i}^{K+1}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}$
(${\leq}$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\frac{\mu_{i}}{2}\frac{\sigma_{i}}{1-\sigma_{i}}\left\|u_{i}^{k+1}-u_{i}\right\|^{2}+\sum_{k=1}^{K-1}\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}}{1-\sigma_{i}}\beta_{i}\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{H_{\beta_{i}}}^{2}$
$+\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2},\quad\forall u\in\mathbb{R}^{mn}.$
This completes the proof. ∎
Now, with the help of Lemma A.1 and Lemma A.2, we can show the convergence of our algorithm in the subsequent proof.
Proof of Theorem 3.3.
First, from (A.5) and (3.12), we can have
[…]
$=$ $\sum_{i=1}^{m}\alpha_{i}\left(u_{i}^{k+1}-u_{i}\right)^{\top}\left(\nabla f_{i}\left(u_{i}\right)-\nabla f_{i}(u_{i}^{k+1})\right)$
(${\leq}$ $-\sum_{i=1}^{m}\alpha_{i}c_{i}\left\|u_{i}-u_{i}^{k+1}\right\|^{2}.$
Then, using (A.11) and (A.23) established in Lemma A.1 and Lemma A.2, respectively, we obtain
$\sum_{k=1}^{K}\left(w^{k+1}-w\right)^{\top}F(w)$
$=$ $\sum_{k=1}^{K}\left(w^{k+1}-w\right)^{\top}F(w^{k+1})+\sum_{k=1}^{K}\left(w^{k+1}-w\right)^{\top}\left(F(w)-F(w^{k+1})\right)$
[…]
$-\sum_{k=1}^{K}\frac{1}{2}\left\|v^{k}-v^{k+1}\right\|_{H_{1}}^{2}-\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}c_{i}\left\|u_{i}-u_{i}^{k+1}\right\|^{2}$
(${\leq}$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\frac{\mu_{i}}{2}\frac{\sigma_{i}}{1-\sigma_{i}}\left\|u_{i}^{k+1}-u_{i}\right\|^{2}-\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}c_{i}\left\|u_{i}-u_{i}^{k+1}\right\|^{2}$
$+\sum_{k=1}^{K-1}\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{{H_{\beta_{i}}}}^{2}-\sum_{k=1}^{K}\sum_{i=1}^{m}\frac{\alpha_{i}}{2}\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{{H_{\beta_{i}}}}^{2}$
$+\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}$
$+\frac{1}{2}\left(\left\|v^{1}-v\right\|_{H_{1}}^{2}-\left\|v^{K+1}-v\right\|_{H_{1}}^{2}\right)$
$=$ $\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\left(\frac{\mu_{i}}{2}\frac{\sigma_{i}}{1-\sigma_{i}}-c_{i}\right)\left\|u_{i}^{k+1}-u_{i}\right\|^{2}+\sum_{k=1}^{K-1}\sum_{i=1}^{m}\frac{\alpha_{i}}{2}\left(\frac{1}{\mu_{i}}\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}-1\right)\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{H_{\beta_{i}}}^{2}$
$+\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}$
$+\frac{1}{2}\left(\left\|v^{1}-v\right\|_{H_{1}}^{2}-\left\|v^{K+1}-v\right\|_{H_{1}}^{2}-\left\|v^{K}-v^{K+1}\right\|_{H_{1}}^{2}\right).$ (A.34)
For the solution point $w^{*}$, it follows from (A.6) that $F(w^{*})=0$. Setting $w=w^{*}$ in (A.1.3), together with the above property, for any integer $K>1$, we have
$\sum_{k=1}^{K}\sum_{i=1}^{m}\alpha_{i}\left(c_{i}-\frac{\mu_{i}}{2}\frac{\sigma_{i}}{1-\sigma_{i}}\right)\left\|u_{i}^{k+1}-u_{i}^{*}\right\|^{2}+\sum_{k=1}^{K-1}\sum_{i=1}^{m}\frac{\alpha_{i}}{2}\left(1-\frac{1}{\mu_{i}}\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}\right)\left\|v_{i}^{k}-v_{i}^{k+1}\right\|_{H_{\beta_{i}}}^{2}$ (A.35)
$\leq$ $\sum_{i=1}^{m}\alpha_{i}\frac{1}{2\mu_{i}}\frac{\sigma_{i}}{1-\sigma_{i}}\left(\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|+\sqrt{\beta_{i}}\left\|v_{i}^{0}-v_{i}^{1}\right\|_{H_{\beta_{i}}}\right)^{2}$
$+\frac{1}{2}\left(\left\|v^{1}-v^{*}\right\|_{H_{1}}^{2}-\left\|v^{K+1}-v^{*}\right\|_{H_{1}}^{2}-\left\|v^{K}-v^{K+1}\right\|_{H_{1}}^{2}\right).$
Note that $\alpha_{i}>0,\forall i\in[m]$, and it follows from (A.10) that
$\left(c_{i}-\frac{\mu_{i}}{2}\frac{\sigma_{i}}{1-\sigma_{i}}\right)>0\quad\text{ and }\quad\left(1-\frac{1}{\mu_{i}}\frac{\sigma_{i}\beta_{i}}{1-\sigma_{i}}\right)>0,\quad\forall i\in[m].$ (A.36)
Therefore, the inequality (A.35) implies that
$\left\|u_{i}^{k+1}-u_{i}^{*}\right\|\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0,\quad\left\|v_{i}^{k+1}-v_{i}^{k}\right\|_{H_{\beta_{i}}}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0,\quad\forall i\in[m].$ (A.37)
Using $\|v_{i}^{k+1}-v_{i}^{k}\|_{H_{\beta_{i}}}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0$ and (A.24), we can infer that
$\left\|z^{k+1}-z^{k}\right\|\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0,\quad\left\|\lambda_{i}^{k+1}-\lambda_{i}^{k}\right\|\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0,\quad\forall i\in[m].$ (A.38)
Since $\|u_{i}^{k+1}-z^{k+1}\|=\frac{1}{\beta_{i}}\|\lambda_{i}^{k+1}-\lambda_{i}^{k}\|$, we can also have
$\left\|u_{i}^{k+1}-z^{k+1}\right\|\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0,\quad\forall i\in[m]$ (A.39)
Combining the facts that $u_{i}^{k}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}u_{i}^{*}$ in (A.37), $u_{i}^{*}=z^{*}$, and $\|u_{i}^{k+1}-z^{k+1}\|\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0$ in (A.39), one has
$z^{k}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}z^{*}.$ (A.40)
It follows from (A.37) that for any $\epsilon>0$, there exists $k_{0}$, such that for all $k\geq k_{0}$, we have ${\|v_{i}^{k+1}-v_{i}^{k}\|_{H_{\beta_{i}}}<\epsilon}$ and $\sigma_{i}^{k}<\epsilon$. Then, for all $k\geq k_{0}$, it follows from (A.30) that
$\left\|e_{i}^{k}(u_{i}^{k+1})\right\|\leq$ $\sum_{j=0}^{k-1}\sigma_{i}^{k-j}\sqrt{\beta_{i}}\|v_{i}^{j}-v_{i}^{j+1}\|_{H_{\beta_{i}}}+\sigma_{i}^{k}\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|$
$=$ $\sum_{j=0}^{k_{0}-1}\sigma_{i}^{k-j}\sqrt{\beta_{i}}\|v_{i}^{j}-v_{i}^{j+1}\|_{H_{\beta_{i}}}+\sum_{j=k_{0}}^{k-1}\sigma_{i}^{k-j}\sqrt{\beta_{i}}\|v_{i}^{j}-v_{i}^{j+1}\|_{H_{\beta_{i}}}+\sigma_{i}^{k}\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|$
$\leq$ $\left(\sqrt{\beta_{i}}\max_{0\leq j\leq k_{0}-1}\|v_{i}^{j}-v_{i}^{j+1}\|_{H_{\beta_{i}}}\sum_{j=0}^{k_{0}-1}\sigma_{i}^{k-k_{0}-j}\right)\sigma_{i}^{k_{0}}+\epsilon\sqrt{\beta_{i}}\sum_{j=k_{0}}^{k-1}\sigma_{i}^{k-j}+\epsilon\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|$
$\leq$ $\epsilon\left(\sqrt{\beta_{i}}\max_{0\leq j\leq k_{0}-1}\|v_{i}^{j}-v_{i}^{j+1}\|_{H_{\beta_{i}}}\sum_{j=0}^{k_{0}-1}\sigma_{i}^{k-k_{0}-j}+\sqrt{\beta_{i}}\sum_{j=k_{0}}^{k-1}\sigma_{i}^{k-j}+\left\|e_{i}^{0}\left(u_{i}^{1}\right)\right\|\right),$
which implies that
$\left\|e_{i}^{k}(u_{i}^{k+1})\right\|\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0,\quad\forall i\in[m].$ (A.41)
From (A.6) and (3.6), we can have $\lambda_{i}^{*}=\nabla f_{i}(u_{i}^{*})$ and
$\lambda_{i}^{k}=\nabla f_{i}(u_{i}^{k+1})+\beta_{i}(u_{i}^{k+1}-z^{k})-e_{i}^{k}(u_{i}^{k+1}).$ (A.42)
Then, we have
$\lambda_{i}^{k}-\lambda_{i}^{*}=\nabla f_{i}(u_{i}^{k+1})-\nabla f_{i}(u_{i}^{*})+\beta_{i}(u_{i}^{k+1}-u_{i}^{k})+\beta_{i}(u_{i}^{k}-z^{k})-e_{i}^{k}(u_{i}^{k+1}).$ (A.43)
Note that $u_{i}^{k}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}u_{i}^{*},u_{i}^{k}-z^{k}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0$, $e_{i}^{k}\left(u^{k+1}\right)\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}0$, and the gradient of $f_{i}$ is Lipschitz continuous (see 3.1), we have
$\lambda_{i}^{k}\stackrel{{\scriptstyle k\rightarrow\infty}}{{\longrightarrow}}\lambda_{i}^{*},\quad\forall i\in[m].$ (A.44)
We thus complete the proof. ∎